### v1

@@ Abstract
Keywords:
Issue resolution Large Language Models GitHub Conversation Analysis
@@ 1 Introduction
There is a growing body of research on using LLMs for issue resolution with a focus on areas like bug localization (Hossain et al, 2024), and program repair (Jimenez et al, 2024; Zhang et al, 2024). Prior research has highlighted the importance of developer conversations in successful issue resolution (Chatterjee et al, 2020; Kavaler et al, 2017). However, detailed analysis of conversations between developers and LLMs for issue resolution remains limited. To address this gap, our previous work (Ehsani et al, 2025) examined 433 developer-ChatGPT conversations in GitHub issue threads, examining how knowledge gaps in prompts and conversation styles affect issue closure. Over 54% of prompts exhibited knowledge gaps, including Missing Context, Missing Specifications, Multiple Contexts, and Unclear Instructions. While we identified critical knowledge gaps in developer prompts, our analysis was limited to understanding what gaps exist and their frequency. We did not explore the broader interaction dynamics, such as how developers adapt to these challenges, how task type, project complexity, or communication patterns contribute to helpful or unhelpful outcomes, or what strategies could mitigate their impact during multi-turn interactions.
Related research has also explored why developers engage in multi-turn conversations with ChatGPT and their motivations for sharing these conversations on GitHub (Hao et al, 2024). A recent work by Das et al. revealed that ChatGPT is primarily used for ideation rather than validation, with generated code rarely being used directly to resolve issues (Das et al, 2025). Mondal et al. identified gaps and factors that prolong these conversations (Mondal et al, 2024). While such conversations do not always lead to successful outcomes, prior studies have not investigated the specific characteristics that make some developer-LLM conversations helpful and others not, or how they contribute to effective issue resolution.
@@
  • We introduce the first manually annotated dataset of helpful and unhelpful developer-ChatGPT issue resolution conversations, consisting of 876 conversations collected from 655 projects, with 686 of the conversations from 543 projects being directly related to issues. Our code and data are made publicly available [^1] https://anonymous.4open.science/r/llm_issue_resolution_analysis-2C53/README.md.
@@
  • We offer actionable strategies for developers to help developers obtain useful ChatGPT suggestions for issue resolution. Our findings can also inform design of prompt optimization tools and fine-tuning LLMs for issue resolution.
@@ 2 Dataset Collection and Preprocessing
In May 2023, OpenAI introduced a feature allowing users to share their conversations with ChatGPT through dedicated links (OpenAI, 2023). Using this feature, Xiao et al. collected a set of developer-ChatGPT conversations publicly shared on GitHub until October 2023, forming the DevGPT dataset (Xiao et al, 2024). The size of this dataset is limited, containing only 413 conversations collected from 2023-01-01 to 2023-10-13. We expand this dataset by collecting additional data using GitHub’s API endpoint, searching for occurrences of ChatGPT links in all GitHub issues. The data is collected from 2023-01-01 to 2024-04-30, encompassing the same window as DevGPT and extending it by more than six months. We gathered all publicly available developer–ChatGPT conversation links, resulting in 876 unique conversation links from 823 GitHub issues across 655 repositories, as shown in Table 2. Each issue thread can contain more than one link to developer-ChatGPT conversations. This dataset is not a sampled subset, but rather the result of an exhaustive crawl of all publicly shared conversation links found in GitHub issues. While prior studies relied on the DevGPT dataset (Ehsani et al, 2025; Mondal et al, 2024), our augmented dataset more than doubles the size and offers significantly broader temporal and project-level coverage. Table 1 provides a side-by-side comparison of the original DevGPT dataset and our extended dataset, highlighting improvements in temporal coverage, issue count, and repository diversity.
@@ 3 Methodology
Direct Implementation: We assessed helpfulness by investigating if ChatGPT’s suggestions were directly applied to resolve issues. We examined whether developers made commits or shared links to pull requests integrating the suggested code or idea, after the link was shared within the issue thread. This assessment involved manual analysis of the pull request or commit content to verify that it aligned with the specific code snippets, patterns, or design ideas proposed in the ChatGPT conversation.
@@
The three types of explicit indicators mentioned above are not mutually exclusive. A single conversation may exhibit one or multiple forms of evidence (e.g., a developer may leave a positive comment and also implement the suggestion). We found 103 instances where code suggestions from ChatGPT lacked any of the three types of explicit indicators. In these cases, we check if the suggested code was directly integrated into the project’s codebase using a code clone detection tool. We extracted subsequent commits made to each project’s git repository after the date that the link was shared in the thread and used the code clone detection tool NiCad (Cordy and Roy, 2011) on each commit. If the tool identified a similarity between the code provided by ChatGPT and the commit made to the project, we mark the conversation as helpful. If no similarity was found, we assume that the conversation was unhelpful for the issue. Among these 103 instances, the distribution of programming languages was as follows: Python (49), JavaScript (15), Java (9), C++ (5), Rust (4), C (2), C# (2), and Swift (1). This distribution resulted in 87 instances written in languages supported by NiCad. The remaining 16 instances involved languages not supported by NiCad, such as YAML, JSON, and TypeScript. For these cases, since no code similarity analysis could be performed and no explicit helpfulness indicators were present in the issue thread, we labeled them as unhelpful. For the rest, we used NiCad. The default threshold for code similarity is set to 30%, which is NiCad’s default setting and has been widely adopted in prior studies for detecting meaningful code clone (Das et al, 2024). This threshold allows for minor edits, such as renaming variables or reformatting, while still capturing structural similarity. Given our goal of identifying cases where developers incorporated ChatGPT’s suggestions, this level of sensitivity was appropriate for our analysis.
@@ 4 Results
  • ∗∗ Indicate differences that are statistically significant with both ($p<0.05$) and ($corrected\ p<0.05$)
@@ 4.2.1 Conversational Metrics
We present the results of Structural and Linguistic metrics in Table 4. These metrics are calculated separately for the developer prompts and ChatGPT responses, as shown in columns (Developers) and (ChatGPT), respectively. For each row in this table, we show the relative difference (Rel. Diff.) by calculating the mean average of the helpful (h1) and unhelpful (h2), and then determining the difference as $(h1-h2)/h1$ between helpful and unhelpful conversations. We bold those that are statistically significant according to the Mann-Whitney U test (p-value<0.05) (Nachar, 2008). The Mann-Whitney U test does not assume normal data distribution and is less affected by outliers and skewed data compared to parametric tests (Hollander et al, 2013), making it suitable for our study. We first discuss the results for the metrics in developer prompts. We also present corrected p-values using the Benjamini-Hochberg (BH) correction (Benjamini and Hochberg, 1995) as an additional reliability measure, to show how many of the differences remain statistically significant after adjustment. We use the BH correction to control the false discovery rate (FDR), reducing the risk of false positives among significant results. With many metrics analyzed, this helps ensure the validity of our findings despite the increased likelihood of false positives from multiple comparisons. In multiple tests, the probability of falsely rejecting at least one null hypothesis (Type I error) increases. This correction adjusts p-values to account for the number of tests, reducing this risk. Unlike stricter corrections like Bonferroni (Armstrong, 2014), which control the family-wise error rate and can be overly conservative, the BH correction controls the FDR. This allows for more discoveries while maintaining a low proportion of false positives among significant results Banerjee et al (2023). Corrected p-values ensure our findings are less likely due to random chance. Of the 17 metrics initially found to be significant, 11 remained statistically significant after adjustment, indicating that these metrics are robust indicators and less likely to be false positives.
@@
ChatGPT Conversations Sentiment #Words #Sentences Median #Prompts Rounds of Conversation
@@ 4.2.2 Project Metrics
  • ∗∗ Indicate differences that are statistically significant with both ($p<0.05$) and ($corrected\ p<0.05$)
@@ 4.2.3 Issue Metrics
  • ∗∗ Indicate differences that are statistically significant with both ($p<0.05$) and ($corrected\ p<0.05$)
@@ 6 Threats to Validity
External Validity. Our findings are based on a dataset where developers shared links to their ChatGPT conversations within GitHub issue threads. Therefore, our results may not generalize to other ChatGPT conversations that developers choose not to share on GitHub. Our observations also might not generalize to developer conversations with other LLMs such as Llama and Copilot. To mitigate these threats, we selected ChatGPT due to its widespread popularity and extensive use among developers. In addition, our extensive dataset, collected from 655 projects comprising 48 programming languages, includes a diverse range of developer-ChatGPT interactions.
@@ 7 Related Work
LLMs for Bug and Issue Resolution. LLMs, especially conversational LLMs, are increasingly used in bug localization, bug resolution, and issue tracking(Hou et al, 2024; Wu et al, 2023; Tang et al, 2023). Previous studies have investigated the effectiveness of LLMs in this context and found that developers often interact with ChatGPT to describe bug symptoms, provide contextual information, and seek guidance on potential solutions (Das et al, 2024). Tao et al. (Tao et al, 2024) developed an LLM-powered multi-agent framework to solve GitHub issue threads, achieving a resolution rate of 13.94%. Q&A platforms like Stack Overflow have been crucial for assisting developers in issue resolution, but their traffic declined with the rise of conversational LLMs (Silva et al, 2024). Several studies compared Q&A forum answers to LLM responses, and found that LLMs struggle with questions related to unpopular frameworks and libraries (Silva et al, 2024). For security questions, ChatGPT responses were correct only 56% of the time (Delile et al, 2023). Despite these concerns, users prefer ChatGPT for their well-articulated language (Kabir et al, 2024). ChatGPT answers were consistently found to be of lower quality than Stack Overflow, often lacking relevance and information timeliness (Xu et al, 2023). Human expertise to solve specific exceptions makes Stack Overflow better in debugging tasks (Liu et al, 2023). Overall, these studies highlight the evolving role of LLMs in issue resolution. We contribute to the existing knowledge by identifying specific SE tasks that LLMs effectively support for issue resolution and evaluating their helpfulness for each task.

### v2

@@ 1 Introduction
There is a growing body of research on using LLMs for issue resolution with a focus on areas like bug localization (Hossain et al, 2024), and program repair (Jimenez et al, 2024; Zhang et al, 2024). Prior research has highlighted the importance of developer conversations in successful issue resolution (Chatterjee et al, 2020; Kavaler et al, 2017). However, detailed analysis of conversations between developers and LLMs for issue resolution remains limited. To address this gap, our previous work (Ehsani et al, 2025b) examined 433 developer-ChatGPT conversations in GitHub issue threads, examining how knowledge gaps in prompts and conversation styles affect issue closure. Over 54% of prompts exhibited knowledge gaps, including Missing Context, Missing Specifications, Multiple Contexts, and Unclear Instructions. While we identified critical knowledge gaps in developer prompts, our analysis was limited to understanding what gaps exist and their frequency. We did not explore the broader interaction dynamics, such as how developers adapt to these challenges, how task type, project complexity, or communication patterns contribute to helpful or unhelpful outcomes, or what strategies could mitigate their impact during multi-turn interactions. Understanding which task and project settings are more likely to lead to helpful outcomes can provide developers with actionable insights for effective issue resolution, helping them both avoid common knowledge gaps and craft better prompts. Related research has also explored why developers engage in multi-turn conversations with ChatGPT and their motivations for sharing these conversations on GitHub (Hao et al, 2024). A recent work by Das et al. revealed that ChatGPT is primarily used for ideation rather than validation, with generated code rarely being used directly to resolve issues (Das et al, 2025). Mondal et al. identified gaps and factors that prolong these conversations (Mondal et al, 2024). While such conversations do not always lead to successful outcomes, prior studies have not investigated the specific characteristics that make some developer-LLM conversations helpful and others not, or how they contribute to effective issue resolution.
@@
  • We introduce the first manually annotated dataset of helpful and unhelpful developer-ChatGPT issue resolution conversations, consisting of 876 conversations collected from 655 projects, with 686 of the conversations from 543 projects being directly related to issues. Our code and data are made publicly available [^1] https://github.com/SOAR-Lab/llm_issue_resolution_analysis.
@@
  • We offer actionable strategies for developers to help developers obtain useful ChatGPT suggestions for issue resolution. Our findings can also inform the design of prompt optimization tools and fine-tuning LLMs for issue resolution.
@@ 2 Dataset Collection and Preprocessing
In May 2023, OpenAI introduced a feature allowing users to share their conversations with ChatGPT through dedicated links (OpenAI, 2023). Using this feature, Xiao et al. collected a set of developer-ChatGPT conversations publicly shared on GitHub until October 2023, forming the DevGPT dataset (Xiao et al, 2024). The size of this dataset is limited, containing only 413 conversations collected from 2023-01-01 to 2023-10-13. We expand this dataset by collecting additional data using GitHub’s API endpoint, searching for occurrences of ChatGPT links in all GitHub issues. The data is collected from 2023-01-01 to 2024-04-30, encompassing the same window as DevGPT and extending it by more than six months. We gathered all publicly available developer–ChatGPT conversation links, resulting in 876 unique conversation links from 823 GitHub issues across 655 repositories, as shown in Table 2. Each issue thread can contain more than one link to developer-ChatGPT conversations. This dataset is not a sampled subset, but rather the result of an exhaustive crawl of all publicly shared conversation links found in GitHub issues. While prior studies relied on the DevGPT dataset (Ehsani et al, 2025b; Mondal et al, 2024), our augmented dataset more than doubles the size and offers significantly broader temporal and project-level coverage. Table 1 provides a side-by-side comparison of the original DevGPT dataset and our extended dataset, highlighting improvements in temporal coverage, issue count, and repository diversity.
@@ 3 Methodology
Direct Implementation: We assessed helpfulness by investigating if ChatGPT’s suggestions were directly applied to resolve issues. We examined whether developers made commits or shared links to pull requests integrating the suggested code or idea, after the link was shared within the issue thread. This process involved a manual, line-by-line analysis of the pull request or commit content, cross-referenced with the specific responses provided by ChatGPT. We looked for clear alignment, including identical or near-identical code snippets, reused logic, distinctive variable or function names, and structural similarities in proposed fixes or strategies. To ensure methodological rigor, the same two authors collaboratively conducted the analysis across all candidate cases. For each instance of potential direct implementation, both authors reviewed the surrounding context (including timestamps, developer comments, and code diffs) to determine whether the changes could be attributed to the ChatGPT suggestions. Any disagreements were resolved through multiple rounds of discussion, during which both authors revisited the evidence, clarified interpretation criteria, and refined the decision logic. Final labels were assigned only after full consensus was reached.
@@
The three types of explicit indicators mentioned above (e.g., textual cues, appreciation emojis, direct implementation) are not mutually exclusive. A single conversation may exhibit one or multiple forms of evidence (e.g., a developer may leave a positive comment and also implement the suggestion). We found 103 instances where code suggestions from ChatGPT lacked any of the three types of explicit indicators. In these cases, we check if the suggested code was directly integrated into the project’s codebase using a code clone detection tool. We extracted subsequent commits made to each project’s Git repository after the date when the link was shared in the thread and used the code clone detection tool NiCad Version 6.2 (Cordy and Roy, 2011) on each commit, comparing it to the code provided by ChatGPT. To extend NiCad’s support for additional languages, we added grammars based on the TXL grammar resources following their documentation [^2] https://txl.ca/txl-resources.html. If the tool identified a similarity between the code provided by ChatGPT and the commit made to the project, we mark the conversation as helpful. If no similarity was found, we assume that the conversation was unhelpful for the issue. Among these 103 instances, 87 instances were written in languages supported by NiCad, either natively or with added grammars and 16 instances of unsupported languages (e.g., YAML, JSON, TypeScript) for which no grammar files were available. The distribution of programming languages was as follows: Python (49), JavaScript (15), Java (9), C++ (5), Rust (4), C (2), C# (2), Swift (1), unsupported (16). The 16 cases in which the language was unsupported were labeled as unhelpful because no code similarity analysis could be performed and no explicit helpfulness indicators were present in the issue thread. For the rest, we used NiCad. The default threshold for code similarity is set to 30%, which is NiCad’s default setting and has been widely adopted in prior studies for detecting meaningful code clone (Das et al, 2024). This threshold allows for minor edits, such as renaming variables or reformatting, while still capturing structural similarity. Given our goal of identifying cases where developers incorporated ChatGPT’s suggestions, this level of sensitivity was appropriate for our analysis.
@@ 4 Results
  • ∗∗ Indicate differences that are statistically significant with both ($p<0.05$) and (correctedp<0.05corrected\ p<0.05)
@@ 4.2.1 Conversational Metrics
We present the results of Structural and Linguistic metrics in Table 4. These metrics are calculated separately for the developer prompts and ChatGPT responses, as shown in columns (Developers) and (ChatGPT), respectively. For each row in this table, we show the relative difference (Rel. Diff.) by calculating the mean average of the helpful (h1) and unhelpful (h2), and then determining the difference as (h1h2)/h1(h1-h2)/h1 between helpful and unhelpful conversations. We bold those that are statistically significant according to the Mann-Whitney U test (p-value<0.05) (Nachar, 2008). The Mann-Whitney U test does not assume normal data distribution and is less affected by outliers and skewed data compared to parametric tests (Hollander et al, 2013), making it suitable for our study. We first discuss the results for the metrics in developer prompts. We also present corrected p-values using the Benjamini-Hochberg (BH) correction (Benjamini and Hochberg, 1995) as an additional reliability measure, to show how many of the differences remain statistically significant after adjustment. We use the BH correction to control the false discovery rate (FDR), reducing the risk of false positives among significant results. With many metrics analyzed, this helps ensure the validity of our findings despite the increased likelihood of false positives from multiple comparisons. In multiple tests, the probability of falsely rejecting at least one null hypothesis (Type I error) increases. This correction adjusts p-values to account for the number of tests, reducing this risk. Unlike stricter corrections like Bonferroni (Armstrong, 2014), which control the family-wise error rate and can be overly conservative, the BH correction controls the FDR. This allows for more discoveries while maintaining a low proportion of false positives among significant results Banerjee et al (2023). Corrected p-values ensure our findings are less likely due to random chance. Of the 17 metrics initially found to be significant, 11 remained statistically significant after adjustment, indicating that these metrics are robust indicators and less likely to be false positives.
@@
ChatGPT
Conversations
Sentiment #Words #Sentences
Median
#Prompts
Rounds of Conversation
@@ 4.2.2 Project Metrics
  • ∗∗ Indicate differences that are statistically significant with both ($p<0.05$) and (correctedp<0.05corrected\ p<0.05)
@@ 4.2.3 Issue Metrics
  • ∗∗ Indicate differences that are statistically significant with both ($p<0.05$) and (correctedp<0.05corrected\ p<0.05)
@@ 6 Threats to Validity
External Validity. Our findings are based on a dataset where developers shared links to their ChatGPT conversations within GitHub issue threads. Therefore, our results may not generalize to other ChatGPT conversations that developers choose not to share on GitHub. Our observations also might not generalize to developer conversations with other LLMs such as Llama and Copilot. To mitigate these threats, we selected ChatGPT due to its widespread popularity and extensive use among developers. In addition, our extensive dataset, collected from 655 projects comprising 48 programming languages, includes a diverse range of developer-ChatGPT interactions. To further assess coverage, we conducted a follow-up crawl for the subsequent 13-month period (May 2024 – May 2025) since the start of our study. We found only 125 additional conversations, compared to the 876 in our original dataset. With 876 out of a known 1,001 total conversations, our sample achieves a ±1.2% margin of error at the 95% confidence level (assuming maximum variability), indicating that our dataset remains highly representative of the available population during the extended timeframe (Cochran, 1977).
@@ 7 Related Work
LLMs for Bug and Issue Resolution. LLMs, especially conversational LLMs, are increasingly used in bug localization, bug resolution, and issue tracking (Hou et al, 2024; Wu et al, 2023; Tang et al, 2023; Ehsani et al, 2025a). Previous studies have investigated the effectiveness of LLMs in this context and found that developers often interact with ChatGPT to describe bug symptoms, provide contextual information, and seek guidance on potential solutions (Das et al, 2024). Tao et al. (Tao et al, 2024) developed an LLM-powered multi-agent framework to solve GitHub issue threads, achieving a resolution rate of 13.94%. Q&A platforms like Stack Overflow have been crucial for assisting developers in issue resolution, but their traffic declined with the rise of conversational LLMs (Silva et al, 2024). Several studies compared Q&A forum answers to LLM responses, and found that LLMs struggle with questions related to unpopular frameworks and libraries (Silva et al, 2024). For security questions, ChatGPT responses were correct only 56% of the time (Delile et al, 2023). Despite these concerns, users prefer ChatGPT for their well-articulated language (Kabir et al, 2024). ChatGPT answers were consistently found to be of lower quality than Stack Overflow, often lacking relevance and information timeliness (Xu et al, 2023). Human expertise to solve specific exceptions makes Stack Overflow better in debugging tasks (Liu et al, 2023). Overall, these studies highlight the evolving role of LLMs in issue resolution. We contribute to the existing knowledge by identifying specific SE tasks that LLMs effectively support for issue resolution and evaluating their helpfulness for each task.